OpenStudy (anonymous):

a rock fliesupward with velocity 26 m/s(52 degree above horizontal).rock rises to maximum height and lands 12m above initial position.(acceleration due to gravity on this planet is 3.7 m/s squared calculate a)max.height b)time rock is in flight c)range of rock

OpenStudy (anonymous):

you had a question about this one?

OpenStudy (anonymous):

yea ..m not getting the max. height right...its 57m....m getting 61m..

OpenStudy (anonymous):

wat i did was vfy at max height is 0 vfy=viy-at 0=26 sin 52-3.4t t=6.03s then h=viyt-1/2g[t ^{2}\] \[h=26 \sin 52(6.03)-\frac{ 3.4 }{ 2 }(6.03)^{2}\] h=61.73 m

OpenStudy (anonymous):

is anything wrong with it???

OpenStudy (anonymous):

a=3.7

OpenStudy (anonymous):

rest is good:)

OpenStudy (anonymous):

i cant believe I made this mistake and asked you.....i m such an idiot...godddd sorryyyyy...

OpenStudy (anonymous):

not at all...! nice work, just a tiny blip :)

OpenStudy (anonymous):

can i ask you a chemistry doubt??

OpenStudy (anonymous):

here's a little quicker way for future reference though \[V _{f}^{2} =V _{i}^{2} -2ah\] \[V _{f}^{2} =0\] \[-V _{i}^{2} = -2ah\] \[\frac{V _{i}^{2} }{ 2a } = h\]

OpenStudy (anonymous):

my chemistry is pretty rusty...

OpenStudy (anonymous):

ohhk....thank u

OpenStudy (anonymous):

mean the equations.....